Find an equation of the line that passes through the point (3, 5) and is perpendicular to the line
8x + 9y − 4 = 0.
(Let x be the independent variable and y be the dependent variable.)

Answer by stanbon(75887)

Find an equation of the line that passes through the point (3, 5) and is perpendicular to the line
8x + 9y − 4 = 0.
---
y = (-8x+4)/9
Slope of that line: -8/9
-----
Slope of perpendicular line: 9/8
---
Form: y = mx + b
Sustitute for y,x,and m; Solve for "b":
5 = (9/8)3 + b
40/8 = 27/8 + b
b = 13/8
-----
Equation:
y = (9/8)x + (13/8)
======================
Cheers,
Stan H.